Let f(x) = 5 + 12x -x^3. Using the differentiation techniques learnt, find




a. The intervals on which f is increasing


1
Expert's answer
2021-12-07T09:22:01-0500

Let "f(x) = 5 + 12x -x^3". Using the differentiation techniques let us find the intervals on which "f" is increasing. Since "f'(x)=12-3x^2=0" implies "x=-2" or "x=2," and "f'(x)>0" for "x\in(-2,2)," we conclude that the function "f" is increasing on the interval "(-2,2)."